# RewardDS: Privacy-Preserving Fine-Tuning for Large Language Models via Reward Driven Data Synthesis

### v1

@@ 6.6 In-depth Analysis of RewardDS Design
Figure 5: Effectiveness of RewardDS design. (a): Performance on medical QA with different privacy budget allocations for generation and reward proxy model training. The allocation of ‘x + (16-x)’ means the privacy budget for training the generation proxy model is set to x, while the reward proxy model is set to (16-x); (b)/(c)/(d): Quality improvement of the synthetic data, on the Medical QA/Financial QA/Code Generation tasks, after applying the Self-Optimizing Refinement module multiple times.
@@
The impact of different privacy budget allocations.
Although the total privacy budget is controlled at $(16,2e^{-5})$, we can allocate more privacy budget to the generation reward model or the reward proxy model. We will investigate the impact of different privacy budget allocations in Figure 5(a). The results indicate that even a small privacy budget for reward model fine-tuning (e.g., “15+1”, with only 1 allocated to the reward model) outperforms the case where no privacy budget is allocated to the reward model (“16+0”). No privacy budget allocated for reward model means that we do not train the reward proxy model on the client for data filtering or refinement. This suggests that even a marginal privacy cost for reward model training can yield substantial benefits. Furthermore, allocating more privacy budget to the reward model will bring only marginal performance improvements.
@@
Impact of Self-Optimizing Refinement on Synthetic Data Quality.
As shown in Alg. 1, we use the self-optimizing refinement module to re-generate synthetic responses and improve quality during each training epoch. To assess the effectiveness of our self-optimizing refinement module (Alg. 1), we track reward scores of synthetic responses across multiple refinement iterations. A higher reward score indicates better synthetic data quality. Figures 5(b-d) demonstrate that synthetic data quality improves gradually through iterative refinement, explaining our method’s superior performance.
@@
Generalizability across different LLM backbones.
We have evaluated our method with different backbones as the target LLM, including Llama-2-7B-chat-hf MetaAI (2023) and Qwen2.5-14B-Instruct Yang et al. (2024b). Table 3 shows that our method maintains superior performance across various backbones, confirming its backbone-agnostic effectiveness. More analysis is in Appendix F.
@@ 7 Conclusion
While RewardDS has demonstrated its effectiveness in medical QA, legal QA, and code generation tasks, it incurs additional training costs for the reward proxy model. Although the model is lightweight, it still requires extra computational resources.
Additionally, due to computational resource constraints, we applied LoRA fine-tuning on the Qwen2.5-14B-Instruct model to validate our method, as discussed in Appendix F. Full-parameter fine-tuning may yield even better performance. Future work will explore larger LLM backbones and additional categories to further demonstrate the effectiveness of our method as computational resources allow.
In the future, we aim to expand our experiments to include more domain-specific tasks and a wider range of LLM backbones. Furthermore, we plan to optimize the local fine-tuning process of the lightweight proxy models on the client side to reduce computational burdens, enhancing the scalability and feasibility of our method.
@@ 8 Ethics Statement
Appendix A Details of Datasets
@@
As Dong et al. (2024) points out, these public datasets suffer from a “data contamination” issue, where some of the data may have been used to train LLMs on the server, causing the models to memorize it and leading to unnaturally high performance. Moreover, the initial datasets are highly redundant, containing many similar samples. To accurately assess the domain-specific performance of different baselines, we should pre-process these datasets. To be specific, firstly, we evaluate the dataset using the Qwen2.5-7B-Instruct model Yang et al. (2024b) and exclude samples with high accuracy, as higher accuracy suggests these samples may have been part of the LLM’s training data and are thus contaminated.
After addressing the contamination issue, we use the Sentence-T5-Base model Ni et al. (2022a) to compute embeddings for each sample and calculate their similarity. This allows us to remove highly similar samples, ensuring deduplication. The pre-processed dataset is then split into private train set, dev set, and test set, with the detailed statistics shown in Table 2. For fair comparison across all methods, we control the size of our sampled synthetic dataset to be twice the size of the private training set, as shown in Table 2.
Appendix B Compared Methods
@@
Vanilla LLM directly transmits the original private datasets to the server and uses Qwen2.5-7B-Instruct model (Yang et al., 2024b) to generate the answer without any privacy protection.
@@
Locally Fine-tuning implements full-parameter fine-tuning (Ding et al., 2023) of the client-side lightweight Qwen2.5-0.5B-Instruct model (Yang et al., 2024b) across individual domain-specific datasets. The optimized model is subsequently used for inference tasks on three benchmark datasets.
@@
As proposed by Kurakin et al. (2023), DP-Generation first uses DP to full-parameter fine-tune Qwen2.5-0.5B-Instruct model as Generation Proxy Model on the client side. Then transmit the Generation Proxy Model to the server for synthetic data sampling. Then, the synthetic data is used to fine-tune the Qwen2.5-7B-Instruct model on the server for inference service.
@@
On the basis of DP-Generation, DP-Instruct (Yu et al., 2024) introduces an additional step to filter the synthetic data. After sampling synthetic data through the Generation Proxy Model, it clusters synthetic instruction datasets using $K$-means clustering on the Sentence-T5-base (Ni et al., 2022b) embeddings. For each real instruction, find the nearest centroid and resample initial synthetic instructions through the privatized histogram. Then, use the resampled synthetic instructions to fine-tune the Qwen2.5-7B-Instruct model on the server.
@@
Proposed by Wang et al. (2024), KnowledgeSG first fine-tune a client-side the Qwen2.5-0.5B-Instruct model $\mathbf{W}_{Loc}$ with DP. Then transmit the model $\mathbf{W}_{DP}$ to the server and generate synthetic data with the model. Next, it filters the synthetic data with BLEU metrics between the synthetic data and original private datasets. The filtered synthetic instructions are fed into the professional model $\mathbf{W}_{Pro}$, which is the Qwen2.5-7B-Instruct in our reproduction, to generate preferable responses corresponding to these instructions. Finally, it uses the generated instructions and responses to fine-tune $\mathbf{W}_{DP}$ and obtain the final desired model for inference.
Appendix C Implementation Details
We use the Qwen2.5-0.5B-Instruct Yang et al. (2024b) as the backbone for both the generation proxy and reward proxy models, and the Qwen2.5-7B-Instruct as the LLM on the server. For DP fine-tuning of the proxy models, we follow the codebase from Li et al. (2024a), training both models for 3 epochs with a batch size of 4 and a gradient accumulation step of 16. We freeze the embedding layer of the backbone and train the other parameters with a learning rate of 4e-5. The privacy budget for fine-tuning both proxy models is set to $(8,1e^{-5})$, leading to a total privacy budget of $(16,2e^{-5})$ due to the sequential composition law of the DP mechanism Abadi et al. (2016). These settings align with established DP deployments such as Apple’s QuickType and Google’s models, as noted by Lukas et al. (2023).
During synthetic data sampling, we use the vLLM framework Kwon et al. (2023) for fast inference, setting the batch size to 32 and sampling 6 candidate responses for each synthetic query. The sampling templates are detailed in Appendix H. For Reward Guided Filtering, we sort the dataset by reward score, split it into $k$ folds, and select the fold with the highest score, setting $k$ to 6 for medical QA, 5 for financial QA, and 8 for code generation. For Self-Optimizing Refinement, we set the number of candidate responses $N$ as 3 for medical QA and code generation, 2 for finanacial QA task. The hyperparameter analysis is provide in Section 6.5 and Appendix E. The generation temperature is 1.0 and top-p is 0.7 to enhance diversity. The templates used for generating feedback are provided in Appendix H.
@@ Appendix C Implementation Details
Appendix D Details of LLM-Judger Training and Evaluation
Since ROUGE-L, ROUGE-1, and PPL metrics do not fully capture the quality of generated outputs in QA tasks, we use the LLM-Judge Zheng et al. (2023) approach to evaluate the generated outputs for medical QA and financial QA tasks.
@@ Appendix D Details of LLM-Judger Training and Evaluation
First, we fine-tune the LLM-Judgers for these domain-specific tasks (medical QA and financial QA). The fine-tuning process is similar to that of our reward proxy model, where we construct preference pair data as training data and use Bradley-Terry loss Liu et al. (2024) for training. The key difference is that we use the more powerful Qwen2.5-13B-Instruct backbone and fine-tune it with the AdamW optimizer, without adding DP noise. We fine-tune the LLM-Judger for 3 epochs with a learning rate of 4e-5.
@@
During evaluation, we provide the LLM-Judger with both the user query and the generated output, allowing the judger to score the outputs. The judge template is provided in Appendix H. We then compare the scores of outputs from our method and other baselines. If the score difference is less than 1, it is considered a tie. Otherwise, the output with the higher score is viewed as the winner.
Appendix E Hyperparameter Analysis
Figure 6: Analysis of hyperparameters including the number of folds $k$ and the number of candidate responses $N$ in Alg. 1 on the financial QA task and code generation task. For financial QA task, to analyze $k$, we set $N=2$ and to analyze $N$, we set $k=5$. For code generation task, to analyze $k$, we set $N=3$ and to analyze $N$, we set $k=8$.
In this section, we conduct additional experiments to analyze the effect of hyperparameters on our methods for the financial QA and code generation tasks.
For the number of partition folds, $k$, it controls the amount of data selected as clean data. As shown in Figure 6, setting $k=5$ yields the best performance for the financial QA task. For the code generation task, as shown in Figure 6, $k=8$ performs best. Larger values of $k$ lead to the exclusion of more synthetic data, which may result in the model overfitting on smaller data subsets and cause performance degradation.
For the number of candidate responses, $N$, a larger $N$ increases the likelihood of selecting better responses from the candidates. However, increasing $N$ also adds more computational cost, and the performance gain is marginal, as illustrated in Figure 6 and Figure 6. Therefore, we set $N=2$ for the financial QA task and $N=3$ for the code generation task.
Appendix F Extension to More LLM Backbones

### v2

@@ 6.5 In-depth Analysis of RewardDS Design
Impact of RewardDS on Synthetic Data Quality and Downstream Performance.
According to Alg.˜1, RewardDS iteratively refines the synthetic data during each training epoch. As shown in Figure 5(a), the reward score of synthetic data gradually increases with iterative refinement, indicating improved data quality.
We also evaluate the downstream performance of the target LLM on the Medical QA task after being fine-tuned on the synthetic data from different refinement stages. The results in Figure 5(b) show that the downstream performance of the fine-tuned LLM also improves significantly with the improvement of the synthetic data quality, strongly highlighting the effectiveness of RewardDS.
Figure 5: Changes of Reward Score and Downstream Performance in RewardDS for the Medical QA Task. Avg RS indicates the average reward score of the synthetic data, while R1, RL, and PPL represent the downstream performance metrics of the target LLM fine-tuned using these synthetic data.
@@
Training Cost Analysis of RewardDS.
As shown in Figure 2, RewardDS introduces additional modules, including Reward Proxy Model Training, Reward-Guided Filtering, and Self-Optimizing Refinement, into the privacy-preserving fine-tuning process of the server-side LLM. We measure the additional time cost of these modules and compare it with that of the original modules: Generation Proxy Model Training and LLM fine-tuning. As shown in Table 2, the additional time cost from our method accounts for only 29.69% of the total time cost, with most of the time consumed by LLM fine-tuning modules. This is primarily due to the use of a lightweight reward proxy model, making the associated modules highly efficient. Overall, the additional time cost introduced by RewardDS is minimal, strongly demonstrating the practicality of our method.
Time (min) Percentage
Initial Modules: 315 70.31%
Generation Proxy Model Training 45 10.04%
LLM fine-tuning 270 60.26%
Additional Modules From RewardDS: 133 29.69%
Reward Proxy Model Training 49 10.94%
Reward Guided Filtering 12 2.67%
Self-Optimizing Refinement 72 16.07%
Table 2: Runtime of different modules in RewardDS for privacy-preserving fine-tuning on Medical QA task. The most time-consuming module is marked in bold.
Moreover, we compare the overall training time and GPU memory usage of RewardDS with other baselines in Table 3. As for training time cost, our method incurs only a small overhead, primarily from dynamically filtering and refining synthetic data to enhance the overall quality of these data. For the GPU memory cost, we only introduce an additional lightweight reward proxy model with 0.5B parameters, which is relatively small and highly efficient. Overall, with just a slight extra cost in training time and GPU memory, our method can achieves significant performance improvements, as shown in Table 1, making it highly practical for real world deployment.
Method Time Cost (min) Total Parameter (B)
DP-Generation 315 7 + 0.5
DP-Instruct 320 7 + 0.5
KnowledgeSG 166 7 + 0.5
RewardDS 448 7 + 0.5 + 0.5
Table 3: Comparison of the time cost and the GPU memory cost between our method and the other baseline methods. The GPU memory cost are measured by the total model parameter amount.
@@
RewardDS vs Filtering according to Reward Score.
According to Figure 5(a), the initial synthetic data contains substantial samples with low reward scores. One straightforward strategy is to filter out these low-quality samples to reduce noise. As shown in Table 4, simply selecting Top 50% synthetic data for fine-tuning can improve the overall data quality and slightly enhance downstream performance. However, filtering also reduces the size of the training set. As more data is discarded, the downstream performance begins to drop due to the limited training data.
In contrast, RewardDS applies Self-Optimizing Refinement to improve the quality of low reward samples instead of discarding them. It can significantly enhance data quality while maintaining a stable dataset size. Consequently, RewardDS achieves superior downstream performance, as demonstrated in Table 4.
Data Count $\uparrow$ Avg RS $\uparrow$ Downstream RL $\uparrow$
Raw Synthetic Data 6420 22.30 10.94
– Select Top 80% 5136 24.05 11.09
– Select Top 50% 3210 25.66 12.43
– Select Top 30% 1926 26.53 11.19
– Select Top 10% 642 27.43 6.82
Refinement by RewardDS 6420 26.82 17.02
Table 4: Comparison between filtering synthetic data only based on Reward Score (RS) and iterative refinement using RewardDS on Medical QA task. Higher Avg RS indicates better overall data quality. The best results are shown in bold, and the second-best results are underlined.
Other Analysis:
Furthermore, we analyze the impact of different privacy budget allocations for our method in Appendix H. Those results in Figure 8 also clearly demonstrate the effectiveness and robustness of our method.
6.6 Hyperparameter Analysis
In this section, we conduct hyperparameter analysis to further prove the effectiveness of our method.
Privacy Budget $\epsilon$.
We evaluate the performance of our method and baseline approaches under varying privacy budgets $\epsilon$. As shown in Figure 6, our method consistently outperforms all baselines, not only under the commonly used setting of $\epsilon=16$, but also under stricter privacy budgets (e.g., $\epsilon=2$, $4$, and $6$). These results demonstrate the superior performance and broad applicability of our method.
Figure 6: Performance of RewardDS and baselines on the Medical QA task under different privacy budgets $\epsilon$ for the Medical QA task. The performance is evaluated using ROUGE-1 (R1) and ROUGE-L (RL) scores.
Synthetic Data Count $L$.
To further validate the effectiveness of our method, we compare its performance with baseline methods under different synthetic data count $L$. As shown in Table 5, RewardDS consistently outperforms all baselines across different data counts, demonstrating its robustness and effectiveness. Notably, even when the synthetic data—for example is extremely scarce, only 1% (642 samples), our method still maintains competitive performance while other baselines suffer from performance degradation. This superiority is mainly attributed to our self-optimizing refinement module, which effectively improves the quality of those limited synthetic data and can maximize their utility in contrast to the baselines.
Synthetic Data Count 642(1%) 1926(30%) 3210(50%) 5136(80%) 6420(100%)
DP-Generation (Kurakin et al., 2023) 6.71 9.55 10.60 10.30 10.94
DP-Instruct (Yu et al., 2024) 6.63 8.00 7.99 8.45 8.44
KnowledgeSG (Wang et al., 2024) 9.43 10.21 10.54 10.92 10.74
RewardDS 16.46 17.17 17.39 17.34 17.02
Table 5: Performance comparison between RewardDS and the other baseline methods under different synthetic data count for the Medical QA task.
More Hyperparameters.
We also analyze more hyperparameters in our method described in Alg.˜1, including the number of folds $k$ and the number of candidate responses $N$. As shown in Figure 9, 10 and 11, our method remains effective and robust across different hyperparameter settings on three domain-specific tasks. More detailed analyses can be found in Appendix I.
6.7 Generalization to More Proxy Models and LLMs
We also evaluate the performance of our method when using different combinations of generation proxy models and reward proxy models, as shown in Table 6. To explore the impact of proxy model heterogeneity, we evaluate the combination of Qwen-0.5B + LLaMA3-1B. To explore the impact of proxy model scale, we consider the combination of Qwen-0.5B + Qwen-1.5B. As shown in the table, using Qwen-0.5B as the generation proxy model and Qwen-1.5B as the reward proxy model achieves the best performance. This is mainly because both proxy models share a similar architecture, which ensures more effective collaboration between them. Moreover, a larger reward proxy model can more effectively enhance the quality of synthetic data and achieve better performance, which underscores the critical role of the reward proxy model.
Generation Proxy Reward Proxy RL Score R1 Score PPL
Qwen-0.5B Qwen-0.5B 17.02 27.78 1.17
Qwen-0.5B Llama3-1B 17.69 28.71 1.21
Qwen-0.5B Qwen-1.5B 18.11 30.30 1.18
Llama3-1B Qwen-0.5B 12.52 19.35 1.15
Qwen-1.5B Qwen-0.5B 17.44 29.64 1.24
Table 6: Performance of our method under different combinations of generation proxy models and reward proxy models for the medical QA task.
Moreover, we evaluate the performance of our method with different server-side LLMs, such as Llama-2-7B-chat-hf MetaAI (2023) and Qwen2.5-14B-Instruct Yang et al. (2024b) in Appendix H. As shown in Table 8, our method consistently outperforms other baselines across various LLMs, demonstrating its effectiveness and robustness.
@@ 7 Conclusion
Due to computational resource constraints, we applied LoRA fine-tuning on the Qwen2.5-14B-Instruct model to validate our method, as discussed in Appendix H.2. Full-parameter fine-tuning may yield even better performance.
Future work will investigate larger LLM backbones to further validate the effectiveness of our method on models with greater parameter scales.
Moreover, although our method is already time-efficient, we plan to further improve efficiency by exploring lightweight training approaches, such as Low-Rank Adaptation (LoRA) and prefix tuning, during the fine-tuning of both the generation proxy model and the reward proxy model.
@@ 8 Ethics Statement
Appendix Overview
The appendix is organized into two parts: Appendix A–D provide related work and the main experimental setup of RewardDS, while Appendix E–J present additional experimental results to further demonstrate the effectiveness of RewardDS.
Appendix A Detailed Related Works
In this section, we provide a more detailed introduction to the related works on LLM privacy-preserving fine-tuning methods, including Anonymity-based methods, Encryption-based methods and Synthesis-based methods.
A.1 Anonymity-based methods
Anonymity-based methods aim to identify and remove user-specific private information from private data to enable privacy-preserving LLM fine-tuning. Sweeney (1997) achieves k-anonymity by dynamically identifying user-specific private information and applying substitution or removal to protect it. Romanov et al. (2019) employs a transformer framework with attention mechanisms to enhance anonymization performance. Staab et al. (2024) proposes an adversarial anonymization approach that leverages one LLM to anonymize user privacy while using another LLM to detect privacy information, iteratively improving the anonymization performance.
All of the above anonymity-based methods require detecting and removing user privacy, which will make the data incoherent and incomplete, thereby reducing its quality for downstream LLM fine-tuning.
A.2 Encryption-based methods
Encryption-based methods focus on applying encryption to the private data and maintaining secure communication between client and server to transmit the private data. Lou et al. (2020) applies fully Homomorphic encryption to protect private data, enabling privacy security while maintaining comparable model performance after fine-tuning on the encrypted data. Frery et al. (2025) combines the Low-Rank Adaptation technique and homomorphic encryption to improve the efficiency of LLM privacy-preserving fine-tuning. You et al. (2025) introduces the hybrid secret sharing algorithm by combining arithmetic secret sharing (ASS) and function secret sharing (FSS) to achieve secure computation during LLM privacy-preserving fine-tuning.
However, current encryption-based methods still require numerous time and resources for encrypting private data and ensuring secure communication, making them impractical for real-world applications.
A.3 Synthesis-based methods
Synthesis-based methods have recently emerged as a more practical and reliable approach, which leverages synthetic data as a substitute for private data in LLM fine-tuning to balance utility and privacy.
Kurakin et al. (2023); Yue et al. (2023) propose using DP-SGD to locally fine-tune a lightweight model on the client as a generation proxy. This proxy model is used to generate synthetic data without privacy risks. The server utilizes the synthetic data to fine-tune the LLM, achieving privacy-preserving training. Considering that those synthetic data often contain numerous incoherent and flawed samples, Yu et al. (2024); Hou et al. (2024) filter out low-quality data by measuring similarity between synthetic and private data. Alternatively, Wang et al. (2024, 2022) avoid sampling synthetic data from the generation proxy model, instead using LLM on the server to improve the proxy model by distillation.
Nevertheless, only text similarity is too surface-level to accurately assess the quality of synthetic data for domain-specific tasks. Moreover, since the server-side LLM is not fine-tuned for domain-specific tasks, its ability to enhance the generation proxy model through distillation is limited.
Appendix B Details of Datasets
@@
As Dong et al. (2024) points out, these public datasets suffer from a "data contamination" issue, where some of the data may have been used to train LLMs on the server, causing the models to memorize it and leading to unnaturally high performance. Moreover, the initial datasets are highly redundant, containing many similar samples. To accurately assess the domain-specific performance of different baselines, we should pre-process these datasets. To be specific, firstly, we evaluate the dataset using the Qwen2.5-7B-Instruct model Yang et al. (2024b) and exclude samples with high accuracy, as higher accuracy suggests these samples may have been part of the LLM’s training data and are thus contaminated.
After addressing the contamination issue, we use the Sentence-T5-Base model Ni et al. (2022) to compute embeddings for each sample and calculate their similarity. This allows us to remove highly similar samples, ensuring deduplication. The pre-processed dataset is then split into private train set, dev set, and test set, with the detailed statistics shown in Table 7. For fair comparison across all methods, we control the size of our sampled synthetic dataset to be twice the size of the private training set, as shown in Table 7.
Appendix C Compared Methods
@@
Vanilla LLM directly uses the LLM (Qwen2.5-7B-Instruct) on the server for those domain-specific tasks.
@@
Locally Fine-tuning fine-tunes a lightweight model (Qwen2.5-0.5B-Instruct) using the private data on the client for those domain-specific tasks.
@@
As proposed by Kurakin et al. (2023), DP-Generation first uses DP to fine-tune a lightweight model (Qwen2.5-0.5B-Instruct) as the Generation Proxy Model on the client side. Then, it transmits the Generation Proxy Model to the server for synthetic data sampling. Then, the synthetic data is used to fine-tune the LLM (Qwen2.5-7B-Instruct) on the server for those domain-specific tasks.
@@
Compared to DP-Generation, DP-Instruct (Yu et al., 2024) introduces a filtering step to improve the quality of synthetic data. After sampling synthetic data from the Generation Proxy Model, it computes the text similarity between the synthetic data and those private data. It filters out those low-similar data to improve data quality. Then the filtered synthetic data is used to fine-tune the LLM (Qwen2.5-7B-Instruct) on the server for those domain-specific tasks.
@@
Considering the high noise in synthetic data, KnowledgeSG Wang et al. (2024) avoids directly sampling synthetic data from the Generation Proxy Model. Instead, it distills knowledge from the LLM to enhance the Generation Proxy Model for domain-specific tasks. Specifically, KnowledgeSG first uses DP to fine-tune a lightweight model (Qwen2.5-0.5B-Instruct) as the Generation Proxy Model on the client. Then, it transmits the Generation Proxy Model to the server and generates synthetic instructions. The synthetic instructions are fed into the professional LLM (Qwen2.5-7B-Instruct) to generate high quality responses. By using the high quality responses to fine-tune the Generation Proxy Model, it can distill the capacity of LLM into the Generation Proxy Model. Finally, the Generation Proxy Model serves for those domain-specific tasks.
Appendix D Implementation Details
We use the Qwen2.5-0.5B-Instruct Yang et al. (2024b) as the backbone for both the generation proxy and reward proxy models, and the Qwen2.5-7B-Instruct as the LLM on the server. For DP fine-tuning of the proxy models, we follow the codebase from Li et al. (2024a), training both models for 3 epochs with a batch size of 4 and a gradient accumulation step of 16. We freeze the embedding layer of the backbone and train the other parameters with a learning rate of 4e-5. The privacy budget for fine-tuning both proxy models is set to $(8,1e^{-5})$, leading to a total privacy budget of $(16,2e^{-5})$ due to the sequential composition law of the DP mechanism Abadi et al. (2016). These settings align with established DP deployments such as Apple’s QuickType and Google’s models, as noted by Lukas et al. (2023). More comparisons between our method and baselines under different privacy budgets are presented in Section 6.6.
During synthetic data sampling, we use the vLLM framework Kwon et al. (2023) for fast inference, setting the batch size to 32 and sampling 6 candidate responses for each synthetic query. The sampling templates are detailed in Appendix J. For Reward Guided Filtering, we sort the dataset by reward score, split it into $k$ folds, and select the fold with the highest score, setting $k$ to 6 for medical QA, 5 for financial QA, and 8 for code generation. For Self-Optimizing Refinement, we set the number of candidate responses $N$ as 3 for medical QA and code generation, 2 for financial QA task. The hyperparameter analysis is provided in Section 6.6 and Appendix I. The generation temperature is 1.0 and top-p is 0.7 to enhance diversity. The templates used for generating feedback are provided in Appendix J.
@@ Appendix D Implementation Details
Appendix E Details of LLM-Judge Evaluation
Considering ROUGE-L/ROUGE-1 metrics only measure lexical similarity to references and PPL only captures fluency, they often fail to assess deeper aspects of response quality. To ensure more reliable evaluation on the generated outputs for the medical QA and financial QA tasks, we adopt the LLM-Judge approach Zheng et al. (2023) for assessment.
@@ Appendix E Details of LLM-Judge Evaluation
First, we fine-tune the LLM-Judger for these domain-specific tasks (medical QA and financial QA). The fine-tuning process is similar to that of our reward proxy model, where we construct preference pair data as training data and use Bradley-Terry loss Liu et al. (2024) for training. The key difference is that we use the more powerful Qwen2.5-13B-Instruct backbone and fine-tune it with the AdamW optimizer, without adding DP noise. We fine-tune the LLM-Judger for 3 epochs with a learning rate of 4e-5.
@@
During evaluation, we provide the LLM-Judger with both the user query and the generated output, allowing the judger to score the outputs. The judge template is provided in Appendix J. We then compare the scores of outputs from our method and other baselines. If the score difference is less than 1, it is considered a tie. Otherwise, the output with the higher score is viewed as the winner.
As shown in Figure 3, our method outperforms other baselines in both the medical QA and financial QA tasks. DP-Generation and KnowledgeSG struggle with noisy samples from synthetic data, leading to poor performance. Although DP-Instruct filters synthetic data by comparing with private data and removing low-similarity samples, it achieves only limited performance gains compared to DP-Generation. This shows that simple similarity measures do not fully capture the quality of synthetic data. Locally Fine-tuning avoids noise from synthetic data by fine-tuning a lightweight proxy model on private data locally, but it still underperforms our method due to the limited learning capacity of the lightweight model for domain-specific knowledge.
Appendix F Case studies
Here, we present a representative example to demonstrate the effectiveness of our method by comparing its generated response with those from baseline methods, including DP-Generation, DP-Instruct, and KnowledgeSG.
As shown in Figure 7, DP-Generation includes repetitive and irrelevant symptoms, such as no facial weakness and no difficulty swallowing, which are not directly related to the user’s query. DP-Instruct avoids repeating unrelated symptoms but still offers unhelpful advice, only suggesting that the user see a doctor without providing any meaningful medical analysis. Similarly, KnowledgeSG offers some advice, like conducting a physical examination, but also fails to provide any professional analysis of the user’s symptoms or potential underlying causes.
In contrast, RewardDS provides a more detailed analysis of the user’s symptoms, offers some possible causes, and suggests feasible advice, such as scheduling a cardiologist appointment and undergoing an ECG test, strongly demonstrating its effectiveness.
Figure 7: A representative example from the medical QA task illustrates the high quality of the response generated by RewardDS, compared to baseline methods. Text highlighted in red indicates meaningless or flawed parts of the answer, while text in green marks meaningful and helpful content. We also provide short analyses explaining the disadvantages or advantages of the generated responses.
Appendix G Details of Privacy Protection Evaluation
In this section, we provide further details about our attack methods, including the Data Extraction Attack (Carlini et al., 2021) and Membership Inference Attack (Yeom et al., 2018; Choquette-Choo et al., 2021). We implement both attacks for our method and the baseline methods and lower attack performance indicates stronger privacy protection capacity.
Data Extraction Attack.
According to Carlini et al. (2021), the Data Extraction Attack aims to recover private fine-tuned data from the fine-tuned model. Specifically, the attackers provide the fine-tuned model with partial prefixes of private data and attempt to reconstruct the corresponding complete private data. We implement this attack on our method and the baselines to evaluate their privacy protection capabilities. The implementation details are as follows:
In our scenario, the private data consists of two components: the user query and corresponding answer, both of which may contain sensitive information. We apply the data extraction attack twice to recover the user query and the answer separately. For the user query, we provide the generation proxy model with the first 10 tokens of the private query and prompt it to reconstruct the complete query. Then, to extract the answer, we provide the model with the previously recovered user query and the first 10 tokens of the private answer, prompting it to generate the full private response. We use greedy decoding during generation and set the maximum output length to 256.
To evaluate the attack’s performance, we utilize the ROUGE-L score between the recovered data (user query and answer) and ground true private data. A higher ROUGE-L score indicates better attack performance.
Membership Inference Attack:
As proposed by Yeom et al. (2018); Choquette-Choo et al. (2021), the membership inference attack aims to determine whether a specific data point was included in the private dataset used for fine-tuning. Specifically, the attackers collect numerous mixed data, which may contain some private data, and utilize the fine-tuned model to judge which one is included in the private data. We implement this attack on both our method and the baselines to assess their privacy protection capabilities. The implementation details are as follows:
To construct the mixed dataset, we apply data augmentation techniques, such as synonym replacement and content rewriting, to the private data and generate synthetic samples that are similar in content but not identical to the original private data. For RewardDS, we input the mixed data into the reward proxy model and obtain the corresponding reward scores. Samples with higher reward scores are considered more likely to be part of the private training data. For the baseline methods (DP-Generation, DP-Instruct, and KnowledgeSG), only the generation proxy model is transferred to the server. We then use this model to compute the Perplexity (PPL) of each sample in the mixed dataset. Samples with lower PPL values are considered as the private data.
To evaluate the effectiveness of the attack, we calculate the F1 score of private data identification. A higher F1 score indicates stronger attack performance and thus weaker privacy protection.
Appendix H More Analysis of RewardDS Design
H.1 The impact of different privacy budget allocations.
As described in Section 6.1, we allocate an equal privacy budget to generation proxy model training and reward proxy model training. To explore the impact of different privacy budget allocations, we vary the privacy budget allocation while keeping the total privacy budget fixed at $(16,2e^{-5})$.
As shown in Figure 8, our method performs consistently well across various allocations, except in the extreme case where no budget is allocated to reward model training (i.e., "16+0"). No budget for reward model means that we do not train the reward proxy model on the client for data filtering or refinement. This phenomenon demonstrates the critical role of the reward model. Notably, even allocating a small budget to the reward model (e.g., "15+1") leads to a significant performance boost over the "16+0" case, suggesting that even a minimal privacy cost for reward model training yields substantial benefits.
Figure 8: Performance on medical QA with different privacy budget allocations for generation proxy model and reward proxy model training. The allocation of ‘x + (16-x)’ means the privacy budget for training the generation proxy model is set to x, while the reward proxy model is set to (16-x);
H.2 Generalizability across more LLM backbones.
As shown in Table 8, RewardDS outperforms other baselines regardless of whether Llama-2-7B-chat-hf or Qwen2.5-14B-Instruct is used as the LLM backbone. This strongly demonstrates that our method is consistently effective, regardless of the LLM backbone. It is worth noting that although Qwen2.5-14B-Instruct has a larger number of parameters compared to Llama-2-7B-chat-hf, our method performs better on the Llama-2-7B-chat-hf model. This is likely due to the use of LoRA fine-tuning on Qwen2.5-14B-Instruct, rather than full-parameter fine-tuning. We believe that applying full-parameter fine-tuning to the Qwen2.5-14B-Instruct model would lead to better performance. Overall, our method consistently achieves superior performance across various LLM backbones, which strongly demonstrates its generalizability.